Hello tutor! Just these three, please!

1)Calculate the standard deviation for the following company based on the forecasts given.

Economic condition

Probability

Return forecast

Expansion

20%

13%

Normal

60%

5%

Recession

20%

-14%

Provide your answer in percent, rounded to two decimals, omitting the % sign.

2) As a financial adviser, you are recommending a well-diversified fund with an expected rate of return of 17% and a standard deviation of 39% to your clients. A client would like to split her investment between your fund and T-bills so that her standard deviation is no more than 27%. What will be her expected return? T-bill's yield 3%. Provide your answer in percent rounded to two digits, omitting the % sign.

3)

Assume that you manage a risky portfolio with an expected rate of return of 15% and a standard deviation of 25%. The T-bill rate is 3% . Suppose your risky portfolio includes the following investments in the given proportions.

Dell

50%

Apple

30%

HP

20%

What are the investment proportions of your client’s overall portfolio, including the position in T-bills, if your client's risk aversion coefficient is 3?

      -       A.       B.       C.       D.   

Dell

      -       A.       B.       C.       D.   

Apple

      -       A.       B.       C.       D.   

HP

      -       A.       B.       C.       D.   

T-Bill

A.

32%

B.

19.2%

C.

12.8%

D.

36%
Answer #1

Since, multiple questions have been posted and each question is independent of another, I have answered the first question.

_____

Question 1:

Step 1: Calculate Expected Return

The expected return is calculated as below:

Expected Return = Probability of Expansion*Return Forecast Under Expansion + Probability of Normal*Return Forecast Under Normal + Probability of Recession*Return Forecast Under Recession

Substituting values in the above formula, we get,

Expected Return = 20%*13% + 60%*5% + 20%*(-14%) = 2.80%

_____

Step 2: Calculate Variance

The value of variance is arrived as below:

Variance = Probability of Expansion*(Return Forecast Under Expansion - Expected Return)^2 + Probability of Normal*(Return Forecast Under Normal - Expected Return)^2 + Probability of Recession*(Return Forecast Under Recession - Expected Return)^2

Substituting values in the above formula, we get,

Variance = 20%*(13% - 2.80%)^2 + 60%*(5% - 2.80%)^2 + 20%*(-14% - 2.80%)^2 = 0.008016

_____

Step 3: Calculate Standard Deviation

The value of standard deviation is determined as follows:

Standard Deviation = (Variance)^(1/2) = (0.008016)^(1/2) = 8.95% (answer)